OpenStudy (anonymous):

Differentiate y=−2tan(cosx).

OpenStudy (anonymous):

Have you tried the chain rule?

OpenStudy (anonymous):

-2 is a constant so you can just leave it as it is

OpenStudy (anonymous):

would it be -2(tan)(-sinx) ?

myininaya (myininaya):

close what is the derivative of tan(x)

OpenStudy (anonymous):

ah so the n-2tan(sinx)+sec^2x

myininaya (myininaya):

chain rule is derivative of outside(inside stays same here) * derivative of inside

OpenStudy (anonymous):

Myininaya :D

OpenStudy (anonymous):

-2sec^2(-sinx)?

myininaya (myininaya):

\[y'=\frac{d}{dx}[-2 \tan(\cos(x))]=-2 \frac{d}{dx}\tan(\cos(x)) \\ =-2 \cdot \frac{d}{dx} \tan(x)|_{\cos(x)} \cdot \frac{d}{dx}\cos(x)\]

OpenStudy (anonymous):

oh wait -2sec^2(cosx)?

myininaya (myininaya):

*the derivative of the inside you almost have it

OpenStudy (anonymous):

-2sec^2(cosx)(-sinx)?

OpenStudy (anonymous):

That looks good! You can simplify it a bit more :)

myininaya (myininaya):

yes as ugly as it is and - * -

OpenStudy (anonymous):

im so bad at math ;;

myininaya (myininaya):

do you see what batman was saying you can simplify ?

OpenStudy (anonymous):

the cos and the sec?

OpenStudy (anonymous):

(2sinx)/(cosx)?

myininaya (myininaya):

no